Complete a common-size vertical analysis in the condensed statement of comprehensive income presented below:

Condensed Statement of Comprehensive Income
Sales revenue
P 480,000

Cost of sales
201,600

Gross margin
278,400

Operating expenses
206,400

Operating income
P 72,000
1
Expert's answer
2021-09-27T11:11:50-0400

Common size analysis express items as percentage of sales as follows

Sales revenue (480000/480000)*100%= 100%

Cost of sales (201600/480000)*100%= 42%

Gross margin(278000/480000)*100%=58%

Operating expenses (206400/480000)=43%

Operating income(72000/480000)*100%=15%